OpenStudy (anonymous):

find the vertex focus and directrix of the parabola. y^2-6y+8x+9=0 The vertex of the parabola is___ (write order pair) The focus is _____ (write order pair) what is the directrix? select the correct choice below. y=__________ OR x=__________

OpenStudy (anonymous):

Where is @directrix lol.

OpenStudy (anonymous):

Complete the square.$$\begin{align*}y^2-6y+8x+9&=y^2-6y+9+8x\\&=(y-3)^2+8x\end{align*}$$Since we have \((y-3)^2+8x=0\) we may rearrange to yield \(-8x=(y-3)^2\). This means our directrix and focus are \(-8/4=-2\) away from the vertex \((0,3)\); since this is a left-opening horizontal parabola (hence the \(-\) and \(y^2\)), meaning our focus is at \((-2,3)\) and our directrix \(x=2\).

OpenStudy (anonymous):